a housewife wishes to mix coffee beans worth Php60 and Php85 a kilogram to obtain 50 kilogram to obtain 50 kilograms of coffee beans worth 70 a kilogram. How many kilogram of each should she use?

Answer by mananth(16946)

price ---------------- quantity
Cofee beans type I60 ----------------xlb
coffee beans type II 85----------------50-x lb
Mixture 70.00 ----------------50

60x+85(50-x)=70 *50
60x+4250-85x=3500
60x-85x =3500-4250
-25x =-750
/ -25
x=30 lb Php60.00 Coffee beans type I
20 lb Php85.00 coffee beans type II